OpenStudy (anonymous):

Please Help Me, If f(x) = 4x – 5 and g(x) = 3 – 2x, find f(x) + g(x). plz help A.2x – 2 B.6x – 2 C.2x – 8 D.6x – 8

OpenStudy (anonymous):

if ur not going to answer this then let me close the question

OpenStudy (anonymous):

f(x) + g(x) = (4x - 5) + (3 - 2x) = 4x - 5 + 3 - 2x = 2x -2